Calculate the d3 tunce between the center of the e
Answer #1

Let the distance is d from earth

Now, for equal force due to earth and moon

G*Mmoon/(3.84 *10^8 - d)^2 = G*Mearth/d^2

7.347 *10^22 /(3.84 *10^8 - d)^2 = 5.98*10^24/d^2

solving for d

d = 3.457 *10^8 m

the distance from earth is 3.457 *10^8 m